The interior angle of a regular polygon with 30 sides is 24 less than each exterior angle of a regular polygon with n sides. Calculate the value of n.
Answer by Alan3354(69443)

The interior angle of a regular polygon with 30 sides is 24 less than each exterior angle of a regular polygon with n sides. Calculate the value of n.
--------
Ext angles of 30-gon = 360/30 = 12 degs.
Exterior angles of the n-gon are 12 + 24 = 36 degs.
360/36 --> n = 10 sides